OpenStudy (anonymous):

At a store kayla bought a watermelon that weighed 54.4 pounds watermelon cost 1.30 a pound there are 16 oz in a pound how much did kayla watermelon cost

Vocaloid (vocaloid):

1 pound = $1.30 54.4 pounds = 54*1.30 = ?

OpenStudy (anonymous):

What about the .4

Vocaloid (vocaloid):

woops, 54.4*1.30

Vocaloid (vocaloid):

good catch

OpenStudy (anonymous):

My answer was 70.72 but my answer choices are 4.42 or 5.23

Vocaloid (vocaloid):

are you sure you didn't mean to type in 54.4 ounces?

OpenStudy (anonymous):

Yeah its ounces

Vocaloid (vocaloid):

yeah, that's what I was thinking, too. that would be a pretty big watermelon! anyway, since there are 16 oz in a pound, then 54.4 ounces = 54.4/16 pounds = 3.4 3.4 pounds * 1.30 = 4.42

OpenStudy (anonymous):

Josh will go to the pool 17 days this month a one day pass to the pool is 4.75 a monthly is 36.50 how much money will josh save by buying a monthly pass

Vocaloid (vocaloid):

monthly cost using the monthly pass = 36.50 monthly cost using the daily pass = 17*4.75 then subtract monthly - daily

Vocaloid (vocaloid):

*daily - monthly

OpenStudy (anonymous):

So it will be 80.75-36.50

Vocaloid (vocaloid):

yes